OpenStudy (jamanorthi):

A rectangular garden is 4 m longer than its wide. If the width is decreased by 1 m, and the length is increases by 5 m, the area is increased by 15m^2. Find the original dimension of the garden.

OpenStudy (anonymous):

l=4+w l*w=A1 (4+w)(w)=A1

OpenStudy (jamanorthi):

i don't get it

OpenStudy (anonymous):

length :l width: w area = l*w l=w+4

OpenStudy (anonymous):

area = l*w=(w+4)(w)

OpenStudy (anonymous):

(w-1)(l+5)=area+15

OpenStudy (jamanorthi):

oh got it.!

OpenStudy (jamanorthi):

thanks